At week 10 of a project below to improve distance learning platform, assume that the planned values (PV) are based on time proportionality but the “% Complete” values for earned values (EV) are from the workers actually doing the tasks.

ACTIVITY

PRE-
DECESSORS

DURATION
(WEEKS)

BUDGET ($)

ACTUAL
COST ($)

%
COMPLETE

A

-

5

500

640

100

B

A

3

300

400

100

C

B

4

320

310

80

D

B

6

600

250

60

E

C,D

3

300

0

0

F

E

5

500

0

0

(a) Find the cost variance (CV), schedule variance (SV), and time variance (TV).

(b) Find the three indexes, namely cost performance index (CPI), schedule performance index (SPI), and cost schedule index (CSI). Is this project favorable or not, please comment.

(c) Find the budget at completion (BAC), estimated cost to complete (ETC), and estimated cost at completion (EAC) for the project.
Answer #1

Answer:

Given that,

At week 10 of a project below to improve distance learning platform, assume that the planned values (PV) are based on time proportionality but the “% Complete” values for earned values (EV) are from the workers actually doing the tasks.

ACTIVITY

PRE-
DECESSORS

DURATION
(WEEKS)

BUDGET ($)

ACTUAL
COST ($)

%
COMPLETE

A

-

5

500

640

100

B

A

3

300

400

100

C

B

4

320

310

80

D

B

6

600

250

60

E

C,D

3

300

0

0

F

E

5

500

0

0

\rightarrow First, I have prepared a Gantt chart based on the above table as shown in the question, Budget is distributed per day and per activity basis.

1 B C D E F G H I J K L
2 Activity Predecessors Duration Budget 1 2 3 4 5 6 7
3 A .... 5 500 100 100 100 100 100
4 B A 3 300 100 100
5 C B 4 320
6 D B 6 600
7 E C, D 3 300
8 F E 5 500
9 Total Budget per day 100 100 100 100 100 100 100
10 Cumulative Budget 100 200 300 400 500 600 700

Formula:

1 B C D E F G H
2 Activity Predecessors Duration Budget 1 2 3
3 A .... 5 500 100 100 100
4 B A 3 300
5 C B 4 320
6 D B 6 600
7 E C, D 3 300
8 F E 5 500
9 Total Budget per day =SUM(F3-F8) =SUM(G3-G8) =SUM(H3-H8)
10 Cumulative Budget =F9 =F10+G9 =G10+H9

(a).

At the end of Day 10,

Total Planned Value = 1160 (Ref- Gantt chart)

Total actual cost=640+400+310+250

=1600

Total earned value=(500 \times 100%)+(300 \times 100%)+(320 \times 80%)+(600 \times 60%)

=1416

Cost Variance= Total earned value- Total actual cost

=1416-1600

=-184(Negative)

Schedule Variance= Total earned value- Total Planned Value

=1416-1160

=256

Time Variance=((10 \times 1416)/1160)-10

=12.207-10

=2.207

(b).

Cost performance Index= Total earned value/Total actual cost

=1416/1600

=0.885

We are getting 88.50 worth of work out of every 100 spent. It is actually bad.

Schedule performance Index= Total earned value/Total Planned Value

=1416/1160

=1.22

We are progressing at 122% of the rate that we have planned, it is actually good.

Cost schedule Index= Cost performance Index \times Schedule performance Index

=0.885 \times 1.22

=1.08

As value is more than 1, hence, it is good for the project.

(c).

Budget at completion or BAC= Total Planned Value for 22 weeks

=2520 (Ref- Gantt chart)

Estimated cost at completion at current CPI=2520/0.885

=2487

Estimated cost to complete at current CPI= Estimated cost at completion at current CPI- Total actual cost

=2487-1600

=887

Estimated cost at completion (considered that current variances are atypical of the future) = Total actual cost +( Budget at completion - Total earned value)

=1600+(2520-1416)

=2704

Estimated cost to complete= Estimated cost at completion- Total actual cost

=2704-1600

=1104